OpenStudy (gokuporter):

Write the equation of the line in point intercept form. Points on the line: (-9,-8) & (2,-4)

Nnesha (nnesha):

i think its `point slope` form or slope intercept form

OpenStudy (gokuporter):

Yeah, I meant to write Slope intercept form.

Nnesha (nnesha):

\[\large\rm y-y_1=m(x-x_1)\] point slope form where m is the slope and (x_1,y_1) represent one of the given point and the slope intercept form is \[\large\rm y=mx+b\] where m is slope and b is y-intercept

Nnesha (nnesha):

do you know the slope formula ? formula to find slope ?

OpenStudy (gokuporter):

Is the slope 4/11?

Nnesha (nnesha):

how did you get that ?? :)

OpenStudy (gokuporter):

I used the slope formula to find the slope.

Nnesha (nnesha):

good whats the formula show your work plz :) what numbers did you substitute in to get 4/11for slope ?

OpenStudy (gokuporter):

\[m=\left[\begin{matrix}y2 - y1 \\ x2 - x1\end{matrix}\right]\]

Nnesha (nnesha):

hmm looks good i guess and its fraction \[\rm \frac{ y_2-y_1 }{ x_2-x_1 }=m\] what did you substitute for x_1,x_2 & y_1,y_2?

OpenStudy (gokuporter):

The points on the line

OpenStudy (gokuporter):

x__-9__1 x___2___2 y___-8___1 y______-4___2 @Nnesha

Nnesha (nnesha):

cool and there is a negative in the formula but remember y's values should be at the top of fraction and x at the bottom and don't forget the sign \[\rm \frac{ -8 -(-4) }{ -9-2 }=m\] now simplify that

OpenStudy (gokuporter):

How is -8 y2?

OpenStudy (gokuporter):

-4/-11?

Nnesha (nnesha):

that's y_1 i thought you typed y_! and x_1 first and then y_2 x_2

Nnesha (nnesha):

\[\rm \frac{ y_1-y_2 }{ x_1-x_2}=\frac{y_2-y_1}{x_2-x_1}\] doesn't matter which one comes first you will get the same answer

Nnesha (nnesha):

and yes that's correct now just cancel out the negative signs and then substitute m for the slope value into the slope intercept form which is y=mx+B and then pick one order (doesn't matter which one u use) substitute x and y for the values in order pair into the equation and then solve for b(y-intercept)

Nnesha (nnesha):

order pair*

OpenStudy (gokuporter):

So 4/11 is positive?

Nnesha (nnesha):

correct.

OpenStudy (gokuporter):

So what do substitute for y and b

Nnesha (nnesha):

alright \[\rm y=mx+B\] m is slope which is 4/11 and b is y-intercept that you have to find so first of all substitute m for 4/11 and pick one of the order pair for x and y substitute x and y for that their values and then solve for b

Nnesha (nnesha):

\[\rm y=\frac{4}{11}x+b\] substitute x and y for (-9,-8) OR for (2,-4) and then solve for b

OpenStudy (gokuporter):

\[-4= \frac{ 4 }{ 11 }2 + b ?\]

OpenStudy (gokuporter):

How do I solve for b?

Nnesha (nnesha):

you should put the parentheses it's m times x so it should be \[\rm -4 =\frac{4}{11}(2)+b\] first multiply slope by the x value

OpenStudy (gokuporter):

8/11?

Nnesha (nnesha):

correct \[-4=\frac{8}{11}+b\] solve for b cancel out the 8/11 from right side

OpenStudy (gokuporter):

Do I multiply it?

Nnesha (nnesha):

it's `8/11 PLUS b` addition so do opposite of addition to cancel it out

OpenStudy (gokuporter):

8/11 - 8/11?

Nnesha (nnesha):

correct subtract -8/11 both sides

OpenStudy (gokuporter):

.-.

Nnesha (nnesha):

wat ?? ,-, subtract 8/11 both sides ?? .-.

Nnesha (nnesha):

i have to go ... can you solve for b ?? do you understand that ? ?

OpenStudy (gokuporter):

No..

Nnesha (nnesha):

alright what part u not able to solve let me know so we can do it together

OpenStudy (gokuporter):

The subtracting from both sides

OpenStudy (gokuporter):

I'm good with everything else.

Nnesha (nnesha):

subtract 8/11 both sides\[\rm -4\color{Red}{-\frac{8}{11}}= \frac{8}{11}\color{Red}{-\frac{8}{11}}+b\] at right side positive 8/11 and negative 8/11 would cancel each other out(That was the goal)

Nnesha (nnesha):

\[\rm -4\color{Red}{-\frac{8}{11}}= \cancel{\frac{8}{11}\color{Red}{-\frac{8}{11}}}+b\] get the common denominator at left side

OpenStudy (gokuporter):

brb

Nnesha (nnesha):

here is an example \[\rm \frac{ a }{ b }-\frac{ c }{ d }\] when the denominators aren't the same multiply the denominators of both fraction  and multiply the numerator of first fraction with the denominator of 2nd fraction multiply the numerator of 2nd fraction with the denominator of first fraction that's it